OpenStudy (anonymous):

Help! Solve the equation with the replacement set {–76, –8, 8, 76} |42 – x | = 34 A.8 B.76 C.–8, 8 D.8, 76

OpenStudy (solomonzelman):

when you have this type equations, you put it the following way. |42 – x | = 34 would form 2 equations. \[42~–~x~ =~34~~~~~~~~~~~~~~~~~~~~~~~~~~~~42~-x~=~-34\]\[Solve~~~for~~~x~~~i n~~~each~~~one,~~~can~~~you~~~do~~~this?\]

OpenStudy (solomonzelman):

Ok, I'll show you, ready?

OpenStudy (solomonzelman):

\[42~-x~-42=34-42~~~->-x=-6~~->x=6\]lets work the second one now.

OpenStudy (solomonzelman):

\[42-x=-34,~~~~42-x-42=-34-42,~~~~-x=-76,~~~~x=76\]

OpenStudy (solomonzelman):

\[Let's~~CHECK!\]

OpenStudy (solomonzelman):

\[|42 – (6) | = 34,~~~->~~~~|34 | =34~~~~~~x=6~~WORKS\]

OpenStudy (solomonzelman):

\[|42 – (72) | = 34,~~~~->~~~~|-34 | =34,~~~~~~x=72~~WORKS\]

OpenStudy (solomonzelman):

\(\huge\color{blue}{Does~~~~this~~~~make~~~~sense?}\)

OpenStudy (anonymous):

so 6, 72 are the answers ?

OpenStudy (solomonzelman):

\(\huge\color{purple}{Yes,~~~~but~~~~do~~~~you~~~see~~why?}\)

OpenStudy (anonymous):

but those are not in the replacement set......

OpenStudy (anonymous):

replacement set {–76, –8, 8, 76}

OpenStudy (solomonzelman):

Yes, I check on the calculator, I didn't add correctly, the process is the smae though. It is 76 and 8.

OpenStudy (solomonzelman):

\(\huge\color{cyan}{Sorry!}\)

OpenStudy (anonymous):

ok :) got a little confused there for a minute but i see how you did that :) thank yo u

OpenStudy (solomonzelman):

\(\huge\color{cornflowerblue}{Anytime!}\)

OpenStudy (solomonzelman):

\[\huge\color{magenta}{Have~~~~a~~~~good~~~~Thanksgiving!}\]